# VPO: Aligning Text-to-Video Generation Models with Prompt Optimization

### v1

@@ 1 Introduction
  • We empirically demonstrate that VPO is a novel and competitive alignment method for text-to-video models, outperforming Diffusion DPO and providing orthogonal benefits.
@@ 3.2 Principle-Based SFT
With a diverse and high-quality set of user queries, we employ an LLM with in-context learning to generate initial optimized prompts. The demonstrations used in this process are carefully crafted to guide the LLM in producing clear, well-structured prompts. Additionally, we instruct the LLM to account for potential safety concerns, ensuring that the generated prompts avoid harmful content, such as depictions of violence or explicit material. The exact prompt template used in this data construction process is provided in the Appendix.
@@ 4.1 Experimental Setup
Our experiments are conducted on several popular video generation models, including the CogVideoX series and Open-Sora 1.2. The CogVideoX series excel in producing coherent, motion-rich, long-duration videos, while Open-Sora 1.2 is an efficient open-source alternative. For CogVideoX, we use the official prompt optimization method based on GLM-4 in-context learning, along with GPT-4o in-context learning as baselines. For Open-Sora 1.2, we adopt the official GPT-4o rewriting method as our baseline. Additionally, we introduce an alternative baseline using the image prompt optimizer, Promptist [10].
@@ 4.4 Safety Evaluation
Method
Align-
ment
Stability
Preser-
vation
Physics Overall
Original Query 1.31 0.29 0.62 0.34 3.77
Iteration 1 1.52 0.31 0.67 0.36 4.15
Iteration 2 1.53 0.31 0.67 0.35 4.17
Iteration 3 1.52 0.32 0.68 0.36 4.18
Iteration 4 1.51 0.31 0.67 0.37 4.17
@@ 7 Prompt Template for Data Construction
Figure 8:
@@
In our experiments, we use LLaMA3-8B-Instruct [7] as the base model to train the prompt optimizer. Both the SFT and DPO stages utilize approximately 10k queries for data construction, including around 1k safety-related queries. For SFT data construction, GPT-4o is used to generate optimized prompts, provide critiques, and refine the optimized prompts. Detailed prompts are provided in Section §7. For SFT training, we set the learning rate to 2e-6 and train for five epochs. The training employs a 0.1 warmup ratio and a batch size of 64. The AdamW optimizer is used with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. In the DPO stage, we sample 4 prompts for each query with a temperature of 0.9. For text-level preference pairs, we also choose GPT-4o to judge and refine. In the DPO stage, we sample four prompts per query with a temperature of 0.9. For text-level preference pairs, GPT-4o is used for judgment and refinement. For video-level preference pairs, we ensure they adhere to text-level principles, selecting prompt pairs with a reward score difference greater than 0.5. This process generates approximately 5k preference pairs for DPO training. The DPO training is performed with a learning rate of 5e-7, $\beta=0.1$, a 0.1 warmup ratio, and a batch size of 16 for one epoch. For both SFT and DPO training, we utilize the DeepSpeed Zero-3 strategy [19]. All experiments are conducted on 8$\times$80G NVIDIA H800 GPUs.

### v2

@@ 1 Introduction
  • We demonstrate that VPO is a novel and competitive alignment method for text-to-video models, outperforming Diffusion DPO and providing orthogonal benefits.
@@ 3.2 Principle-Based SFT
With a diverse and high-quality set of user queries, we employ an LLM, specifically GPT-4o, with in-context learning to generate initial optimized prompts. The demonstrations used in this process are carefully crafted to guide the LLM in producing clear, well-structured prompts. Additionally, we instruct the LLM to account for potential safety concerns, ensuring that the generated prompts avoid harmful content, such as depictions of violence or explicit material. The exact prompt template used in this data construction process is provided in the Appendix.
@@ 4.1 Experimental Setup
Our experiments are conducted on several popular video generation models, including the CogVideoX series and Open-Sora 1.2. The CogVideoX series excel in producing coherent, motion-rich, long-duration videos, while Open-Sora 1.2 is an efficient open-source alternative. For CogVideoX, we use the official prompt optimization method based on GLM-4 in-context learning, along with GPT-4o in-context learning as baselines. For Open-Sora 1.2, we adopt the official GPT-4o rewriting method as our baseline. Additionally, we introduce an alternative baseline using the image prompt optimizer, Promptist [10]. Moreover, we compare VPO with Prompt-A-Video [13].
@@ 6 Conclusion
7 Acknowledgement
This work was supported by the National Science Foundation for Distinguished Young Scholars (with No. 62125604). This work was also supported by Tsinghua University Initiative Scientific Research Program. We would also like to thank Zhipu AI for sponsoring GPU computing and API cost consumed in this study.
@@ 8 Prompt Template for Data Construction
Figure 8: Prompt template for critique and refinement generation.
@@
In our experiments, we use LLaMA3-8B-Instruct [7] as the base model to train the prompt optimizer. Both the SFT and DPO stages utilize approximately 10k queries for data construction, including around 1k safety-related queries. For SFT data construction, GPT-4o is used to generate optimized prompts, provide critiques, and refine the optimized prompts. Detailed prompts are provided in Section §8. For SFT training, we set the learning rate to 2e-6 and train for five epochs. The training employs a 0.1 warmup ratio and a batch size of 64. The AdamW optimizer is used with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. In the DPO stage, we sample 4 prompts for each query with a temperature of 0.9. For text-level preference pairs, we also choose GPT-4o to judge and refine. In the DPO stage, we sample four prompts per query with a temperature of 0.9. For text-level preference pairs, GPT-4o is used for judgment and refinement. For video-level preference pairs, we ensure they adhere to text-level principles, selecting prompt pairs with a reward score difference greater than 0.5. This process generates approximately 5k preference pairs for DPO training. The DPO training is performed with a learning rate of 5e-7, $\beta=0.1$, a 0.1 warmup ratio, and a batch size of 16 for one epoch. For both SFT and DPO training, we utilize the DeepSpeed Zero-3 strategy [19]. All experiments are conducted on $\times8$0G NVIDIA H800 GPUs.
@@
10 Comparison with VBench Long Prompts
We show the comparison with VBench Long Prompts in Table 5. VPO consistently outperforms the VBench Long Prompts baseline.
Method
Human
Action
Scene
Multiple
Objects
Appear.
Style
VBench Long Prompts (2B) 98.00 51.33 63.81 24.07
VPO (2B) 99.00 55.83 70.17 24.20
VBench Long Prompts (5B) 98.40 53.67 65.67 24.41
VPO (5B) 99.60 55.68 75.73 24.57
Table 5: Comparison with VBench Long Prompts.
Method
Align-
ment
Stability
Preser-
vation
Physics Overall
Original Query 1.31 0.29 0.62 0.34 3.77
Iteration 1 1.52 0.31 0.67 0.36 4.15
Iteration 2 1.53 0.31 0.67 0.35 4.17
Iteration 3 1.52 0.32 0.68 0.36 4.18
Iteration 4 1.51 0.31 0.67 0.37 4.17
11 Iterative Improvement